OpenStudy (anonymous):

–5(1 – 5x) + 5(–8x – 2) = –4x – 8x A. x = –5 B. x = 0.19 C. x = 0.55 D. x = 5

OpenStudy (anonymous):

plug i A and you'll see it is correct

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

-5(1-5x)+5(-8x-2)=-4x-8x -5+25x-40x-10=-12x -15x-15=-12x -3x=15 x=-5